# Learning Safe Strategies for Value Maximizing Buyers in Uniform Price Auctions

### v2

@@ Appendix E Simulation Details
Dataset. We estimate the richness ratio for the EU ETS emission permit auction data for 2022 and 2023. However, only aggregate statistics of the submitted bids is publicly available for privacy reasons. Hence, we synthesize individual level bid data from these available statistics. The exact procedure to reconstruct the bid data is presented in Section E.1. We sample the values from the $\text{Unif}[0,1]$ distribution. In each simulation, we sample $T\sim\text{Unif}[100,300]$ auctions and let $M\sim\text{Unif}[10,80]$. We vary $m=1$ to $m=10$ and average over 100 simulations to obtain plots in Fig. 4. As computing $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ (the value obtained by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$) can be non-trivial, we obtain a uniform upper bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ that is independent of $m$ (see details in Section E.2).
E.1.  Reconstructing Individual Bid Data
@@
We obtained the publicly available auction data for $T_{\max}=443$ EU ETS emission permit auctions held in 2022 and 2023 (EEX, 2023). For each auction indexed by $t\in[T_{\max}]$, we have the following relevant information: the minimum bid ($b_{\min}^{t}$), the maximum bid ($b_{\max}^{t}$), the average of the bids ($b_{\text{avg}}^{t}$), the median of the bids ($b_{\text{med}}^{t}$), and the number of bid-quantity pairs submitted ($n_{\text{sub}}^{t}$). We normalized the bids to be in $[0,1]$. For all rounds $t$, $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t}$ (linear regression yields coefficient 1.01 and intercept $-0.008$).
@@
Upon further investigation, we observed that, except a few, a significant number of auctions had either $b_{\min}^{t}\approx b_{\text{avg}}^{t}\ll b_{\max}^{t}$ (Type I) or $b_{\min}^{t}\ll b_{\text{avg}}^{t}\approx b_{\max}^{t}$ (Type II). As $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t},\forall t$, we deduce that for Type I, most of the bids are concentrated in the interval $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ whereas for Type II, most of the bids are in the interval $[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$. We posit that for Type I (resp. Type II) auctions, $f\in(0,1)$ fraction of the bids ($n_{\text{sub}}^{t}$) are in $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ (resp. $[2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$) and the $1-f$ fraction of bids are in $[2b_{\text{avg}}^{t}-b_{\min}^{t},b_{\max}^{t}]$ (resp. $[b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\max}^{t}]$). If for Type I (resp. Type II) auctions, $2b_{\text{avg}}^{t}>b_{\min}^{t}+b_{\max}^{t}$ (resp. $2b_{\text{avg}}^{t}<b_{\min}^{t}+b_{\max}^{t}$), we assume that all the bids are uniformly present in the interval $[b_{\min}^{t},b_{\max}^{t}]$. With these assumptions, we generate individual bid data for each auction by sampling uniformly from these intervals.
@@
E.2.  An Uniform Upper Bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right).$
@@
Appendix F Resolving Ties
We first emphasize that the effect of resolution of ties on the objective function under the value maximization behavioral model is fundamentally different from the quasilinear utility maximization model. To illustrate this, consider a single-item second price auction with two bidders each valuing the item equally at $v$. Assume that both the bidders bid $v$ and there exists a definitive tie breaking rule (either randomized or deterministic). If the bidders are considered quasilinear utility maximizers, the objective function value (value obtained minus the payments) for both the bidders is 0. However, under the value maximization model (assume $\gamma=0$ for both the bidders), the objective function value for the winning bidder is $v$ and 0 for the losing bidder. [^14] Although, the total (liquid) welfare is $v$ in both the cases. Hence, we need to carefully analyze the key results of our work in the event of ties.
In the context of multi-unit auctions, we can classify ties into two informal types: (a) ‘good ties’—ties occurring at any bid other than the last accepted bid (LAB) and (b) ‘bad ties’, which are the ties occurring at the LAB. All the results in this work are unaffected in case only ‘good ties’ occur. Thus, we focus on ‘bad ties’ for the rest of this section. For this discussion, we consider a public, deterministic tie breaking rule under which ties are always broken in favor of lower indexed bidder. In the presence of ties, the undominated class of safe bidding strategies is still the class of nested strategies (Theorem 3.2). The tie breaking rule is incorporated into the decomposition in Lemma 4.1 and computing the edge weights in the DAG in the offline and online settings in Section 4, ensuring that the maximum weight path in the DAG gives the optimal offline solution and Algorithm 1 achieves sublinear regret in the online setting.

### v4

@@ Appendix F Simulation Details
We sample the values from the $\text{Unif}[0,1]$ distribution. In each simulation, we sample $T\sim\text{Unif}[100,300]$ auctions and let $M\sim\text{Unif}[10,80]$. We vary $m=1$ to $m=10$ and average over 100 simulations to obtain plots in Fig. 4. As computing $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ (the value obtained by the optimal bidding strategy in $\mathscr{F}^{\mathcal{H}_{-}}_{m}$) can be non-trivial, we obtain a uniform upper bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right)$ that is independent of $m$ (see details in Section F.2).
F.1. Reconstructing Individual Bid Data
@@
We obtained the publicly available auction data for $T_{\max}=443$ EU ETS emission permit auctions held in 2022 and 2023 (EEX, 2023). For each auction indexed by $t\in[T_{\max}]$, we have the following relevant information: the minimum bid ($b_{\min}^{t}$), the maximum bid ($b_{\max}^{t}$), the average of the bids ($b_{\text{avg}}^{t}$), the median of the bids ($b_{\text{med}}^{t}$), and the number of bid-quantity pairs submitted ($n_{\text{sub}}^{t}$). We normalized the bids to be in $[0,1]$. For all rounds $t$, $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t}$ (linear regression yields coefficient 1.01 and intercept 0.008-0.008).
@@
Upon further investigation, we observed that, except a few, a significant number of auctions had either $b_{\min}^{t}\approx b_{\text{avg}}^{t}\ll b_{\max}^{t}$ (Type I) or $b_{\min}^{t}\ll b_{\text{avg}}^{t}\approx b_{\max}^{t}$ (Type II). As $b_{\text{avg}}^{t}\approx b_{\text{med}}^{t},\forall t$, we deduce that for Type I, most of the bids are concentrated in the interval $b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ whereas for Type II, most of the bids are in the interval $2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$. We posit that for Type I (resp. Type II) auctions, $f\in(0,1)$ fraction of the bids ($n_{\text{sub}}^{t}$) are in $b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\min}^{t}]$ (resp. $2b_{\text{avg}}^{t}-b_{\max}^{t},b_{\max}^{t}]$) and the 1f1-f fraction of bids are in $2b_{\text{avg}}^{t}-b_{\min}^{t},b_{\max}^{t}]$ (resp. $b_{\min}^{t},2b_{\text{avg}}^{t}-b_{\max}^{t}]$). If for Type I (resp. Type II) auctions, $2b_{\text{avg}}^{t}>b_{\min}^{t}+b_{\max}^{t}$ (resp. $2b_{\text{avg}}^{t}<b_{\min}^{t}+b_{\max}^{t}$), we assume that all the bids are uniformly present in the interval $b_{\min}^{t},b_{\max}^{t}]$. With these assumptions, we generate individual bid data for each auction by sampling uniformly from these intervals.
@@
F.2. An Uniform Upper Bound for $V^{\mathsf{OPT}}_{m}\left(\mathcal{H}_{-}\right).$
@@
Appendix G Resolving Ties
We first emphasize that the effect of resolution of ties on the objective function under the value maximization behavioral model is fundamentally different from the quasilinear utility maximization model. To illustrate this, consider a second price auction with a single indivisible item with two bidders each valuing the item equally at $v$. Assume that both the bidders bid $v$ and there exists a definitive tie breaking rule (either randomized or deterministic). If the bidders are considered quasilinear utility maximizers, the objective function value (value obtained minus the payments) for both the bidders is 0. However, under the value maximization model (assume $\gamma=0$ for both the bidders), the objective function value for the winning bidder is $v$ and 0 for the losing bidder. [^25] Although, the total (liquid) welfare is $v$ in both the cases. So, we need to carefully analyze the key results of our work in the event of ties.
In the context of multi-unit auctions, we can classify ties into two informal types: (a) ‘good ties’—ties occurring at any bid other than the last accepted bid (LAB) and (b) ‘bad ties’, which are the ties occurring at the LAB. All the results in this work are unaffected in case only ‘good ties’ occur. Thus, we focus on ‘bad ties’ in this paper. For this discussion, we consider a public, deterministic tie breaking rule under which ties are always broken in favor of lower indexed bidder. In the presence of ties, the undominated class of safe bidding strategies is still the class of nested strategies (Theorem 3.2). The tie breaking rule is incorporated into the decomposition in Lemma 4.1 and computing the edge weights in the DAG in the offline and online settings in Section 4, ensuring that the maximum weight path in the DAG gives the optimal offline solution and Algorithm 1 achieves sublinear regret in the online setting.